when 20.0 grams of calcium reacts with hydrochloric acid, hydrogen gas is collected at 350 kelvin and 2.5 atmospheres. what is the volume of the hydrogen gas
1
Expert's answer
2019-06-04T05:35:58-0400

2.5 atm = 253313 Pa


The following reaction occurs:

"Ca \ + \ 2HCl \ = \ CaCl_2 \ + \ H_2\uparrow"

First calculating amount of calcium given:

"n(Ca)=\frac{m(Ca)}{M(Ca)}=\frac{20.0g}{40\frac{g}{mol}}=0.5mol"

The same amount of hydrogen gas will be evolved: n(H2)=n(Ca)=0.5 mol

Using ideal gas law pV=nRT we can find volume of the hydrogen:

"V(H_2)=\frac{n(H_2)RT}{p(H_2)}=\frac{0.5mol \cdot 8.314 \frac{J}{mol \cdot K} \cdot 350K}{253313Pa}=0.00574 m^3=5.74L"